OpenStudy (anonymous):

5 - 4x < -7 a. Solve the inequality showing all steps and b. Describe the graph. (Guide: open or closed circle on what number, arrow to the right or left.)

OpenStudy (igreen):

First step, subtract 5 to both sides..what do you get?

OpenStudy (anonymous):

5-5-4x<-7-5 -4x<-12

OpenStudy (igreen):

Correct, now divide -4 to both sides.

OpenStudy (anonymous):

Answer x>3 I think that is correct

OpenStudy (igreen):

@Thehulk49 Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. Please read the \(\bf\color{blue}{\underline{{\href{http:///openstudy.com/code-of-conduct}{Code-of-Conduct}}}}\) to learn more.

OpenStudy (anonymous):

Answer x>3 I think that is correct

OpenStudy (anonymous):

-4x/-4x<-12/-4x x=3

OpenStudy (anonymous):

its ok i knew the answer i just needed the formula

OpenStudy (igreen):

Lol, x doesn't equal 3..

OpenStudy (igreen):

We're dealing with inequalities.

OpenStudy (anonymous):

oh i know i just accidently put an = sign force of habit

OpenStudy (igreen):

Okay, so what is our final answer?

OpenStudy (anonymous):

x>3

OpenStudy (anonymous):

its because we divided that the sign swithed around right?

OpenStudy (igreen):

Yep, you got it.

OpenStudy (igreen):

Divided by a negative number.

OpenStudy (igreen):

Okay, if it's < or > it's an open dot, if it's \(\le\) or \(\ge\) it's a closed dot. If it's < or \(\le\) it points to the left, if it's > or \(\ge\) it points to the right.

OpenStudy (anonymous):

yea i knew all of that and this i just needed the order to solve the inequality in

OpenStudy (anonymous):

bruh i just got two comments deleted in openstudy feedback in a row

OpenStudy (anonymous):

and i was issued a warning

OpenStudy (anonymous):

first comment was fan all moderators so i know when their around then spam when their gone

OpenStudy (anonymous):

the second was they deleted my comment